Consider the sample: 9, 6, 7, 9, 10, 8

a) Find : the mean, the median, the mode

b) Find : the range, the quartiles Q1, Q3, the variance, the standard deviation.

c) Draw a box-plot for the given data set.
Answer #1

a ) Mean :  

6 49 6 1. =8.1667 = = 6

Median :

The observations need to be arranged in ascending order first :

6 , 7 , 8 , 9 , 9 , 10

We have 6 observations . The median is the average of the third and fourth observation .

Median = ( 8+9)/2=8.5

Mode = 9 ( the observation having the highest frequency )

b ) Range = 10 - 6 = 4

First quartile =  7.25

Third quartile = 9

Variance :

(x, -T)2 - 2.1667 rl

Standard deviation :

7L Σ(r,-z)2 = 1.4720 rl

c ) The box plot is given below :